Identify the equations and illustrate how to calculate the distance travelled by an object, if the only data available was the time travelled, the final velocity, the mass of the object and the force exerted on the object.
1
Expert's answer
2011-10-18T12:18:37-0400
V0=0
S0=0
Needed relations:
According to Newton's second law a=Fr/m
V=V0+at
S=S0+V0t+at^2/2

1. t

S=at^2/2

2. V

a=V/t
S=Vt^2/2t=Vt/2

3. m,F

a=F/m
S=Ft^2/2m